OpenStudy (anonymous):

integral of (sec^2x)(cosec^2x)

OpenStudy (anonymous):

\[ \frac{\sec ^2(x)}{\csc ^2(x)}=\frac{1}{\frac{\cos ^2(x)}{\sin ^2(x)}}=\tan ^2(x) \]

OpenStudy (aravindg):

Isn't it \((\sec^2 x)(cosec^2x)\)?

OpenStudy (anonymous):

\[ \int \tan ^2(x) \, dx=\int \left(\sec ^2(x)-1\right) \, dx=\tan (x)-x +C \]